OpenStudy (anonymous):

What is the first derivative of A(t)= 2000(1.00375)^12t??

OpenStudy (anonymous):

d/dt(2000 1.00375^(12 t)) | Factor out constants: = | 2000 (d/dt(1.00375^(12 t))) | Use the chain rule, d/dt(1.00375^(12 t)) = ( d1.00375^u)/( du) ( du)/( dt), where u = 12 t and ( d1.00375^u)/( du) = 0.00374299 1.00375^u: = | 2000 (0.00374299 1.00375^(12 t) (d/dt(12 t))) | Factor out constants: = | 7.48597 1.00375^(12 t) (12 (d/dt(t))) | The derivative of t is 1: = | 89.8317 1 1.00375^(12 t)